Run summary for commissioning in spring 2016

From Hall A Wiki
Jump to: navigation, search

BPM calibration runs (Bull's eye scan)

Relative position LHRS CODA run RHRS CODA run Harp scan entry no.
(0,0) 12454 21781 3379157
(2,2) 12455 21782 3379165
(2,-2) 12456 21783 3379172
(-2,-2) 12457 21784 3379184
(-2,2) 12458 21785 3379191
(0,0) 12459 21786 3379201

BCM calibration runs

LHRS CODA run RHRS CODA run BCM logger Comments
12382 21731 /a/opsuser/hacuser/BCMLOG/bcm_log_02-07-2016-23-24-07.log
12514 21807 /a/opsuser/hacuser/BCMLOG/bcm_log_02-16-2016-20-07-42.log The details can be found in halog entry 3380376 and comment therein
12916 22097 /a/opsuser/hacuser/BCMLOG/bcm_log_03-04-2016-09-40-14.log Max 30uA (see strip chart); comments and halog entry about beam loss
13220 22298 /a/opsuser/hacuser/BCMLOG/bcm_log_04-10-2016-12-52-33.log See BCM calibration
13447 22476 /a/opsuser/hacuser/BCMLOG/bcm_log_04-21-2016-11-24-33.log See BCM calibration and Another BCM data log

Elastic production runs

  • 2nd Pass (Ebeam = 4.5 GeV)
Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Q2 [GeV2] Comment
21799 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Only one set of raster coils is in use; ps1,2,3=1, ps8=100
21809 - 21820 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Only one set of raster coils is in use; ps1,2,3=1, ps8=100; Signals for 16 straws missing in all three V planes
21824 - 21831 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Same as above
21832 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Beam tripped in the middle of this run, production only lasted for an hour (run time is about 3 hours); same as above
21844 - 21857 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Straw chamber issues in V2 plane still exist
21869 - 21871 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Same as above
21872 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Beam turned off 30 mins after run started; same as above
21875 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Asked MCC to restore beam position 25 mins after run started; Straw chamber issues in V2 plane still exist
21876 - 21887 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Straw chamber issues in V2 plane still exist
21889 - 21893 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Straw chamber issues in V2 plane still exist
21894 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Beam position moving second half of run; straw chamber issues in V2 plane still exist
21895 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Last 16 minutes no beam; straw chamber issues in V2 plane still exist
21896 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Straw chamber issues in V2 plane still exist
21897 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Last 25 minutes no beam; straw chamber issues in V2 plane still exist
21898 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Straw chamber issues in V2 plane still exist
21900 - 21901 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Straw chamber issues in V2 plane still exist
21903 - 21915 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Straw chamber issues in V2 plane still exist
21918 - 21925 15cm LH2 2 x 2 mm2 4.481 1.547 52.91 5.5 Straw chamber issues in V2 plane still exist
  • 4th Pass (Ebeam = 8.8 GeV)
Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Q2 [GeV2] Comment
22027 - 22034, 22036 - 22044, 22047 - 22055, 22058 - 22062 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200
22067 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 Very short run; ps1,2,3=1, ps8=200
22076 - 22077 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200
22079 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 Very short run; ps1,2,3=1, ps8=200
22080 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 Very short run; 10uA; ps1,2,3=1, ps8=200
22081 - 22083 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 Beam very trippy; 10uA; ps1,2,3=1, ps8=200
22085 - 22087 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 10uA; ps1,2,3=1, ps8=200
22088 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 Very short run; 10uA; ps1,2,3=1, ps8=200
22093 - 22095 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200
22098 - 22102 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200
22103 - 22108 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200
22114 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 No beam during last 25 mins; ps1,2,3=1, ps8=200
22115 - 22128 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200
22130 - 22154 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200; beam off for much of run 22146
22155 - 22157 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200; 5 uA
22158 - 22160 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200
22161 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200; very short run
22162-22167 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200; these are all short runs (about 30-40 mins)
22168-22172 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200
22174-22181 15cm LH2 2 x 2 mm2 8.841 2.101 48.75 12.65 ps1,2,3=1, ps8=200
  • Still 4 pass here, but spectrometer moved for DVCS calibration and then moved back before the following runs.
  • No survey was done at this position, but a point run 22297 was taken on multi-foil carbon target at this position.
  • Raster size in MCC unit also changed (check image of carbon hole target and raster current in start- and end-of-run for details)
Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Q2 [GeV2] Comment
22263 - 22264, 22266 - 22293 15cm LH2 2 x 2 mm2 8.841 2.101 48.89 12.65 ps1,2,3=1, ps8=200
22300 - 22301, 22303 - 22314 15cm LH2 2 x 2 mm2 8.841 2.101 48.89 12.65 ps1,2,3=1, ps8=200
  • 5th Pass (Ebeam = 11.0 GeV)
Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Q2 [GeV2] Comment
22196-22197,22202-22207, 22209-22212 15cm LH2 2 x 2 mm2 11.021 2.204 48.75 16.54 ps1,2,3=1, ps8=200
22225-22229, 22235-22236, 22238, 22240-22241 15cm LH2 2 x 2 mm2 11.021 2.204 48.75 16.54 ps1,2,3=1, ps8=200
  • Spectrometer moved for DVCS calibration after 4-pass run 22314. No survey was done at this position, but a point run 22328 was taken on multi-foil carbon target at this position.
Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Q2 [GeV2] Comment
22326, 22329 - 22330, 22333 - 22342, 22346 - 22349, 22352 - 22361 15cm LH2 2 x 2 mm2 11.021 2.204 48.75 16.54 ps1,2,3=1, ps8=200
22366 - 22370, 22379 - 22406 15cm LH2 2 x 2 mm2 11.021 2.204 48.75 16.54 ps1,2,3=1, ps8=200; Raster failed and repaired between run 22361 and 22366, look at this link and here to see what happened
22407 - 22409 15cm LH2 2 x 2 mm2 11.021 2.204 48.75 16.54 ps1,2,3=1, ps8=200; trippy beam
22410 - 22413, 22415 15cm LH2 2 x 2 mm2 11.021 2.204 48.75 16.54 ps1,2,3=1, ps8=200
22419 - 22426, 22428 - 22429 15cm LH2 2 x 2 mm2 11.021 2.204 48.75 16.54 ps1,2,3=1, ps8=200; raster failed and fixed before run 22419; raster size was set to 2(H)x2(W) in MCC units (it was 2(H)X3(W) in MCC units before the raster failure); run 22427 was during Compton setup
22430, 22432 - 22438, 22440 - 22452, 22456 - 22459 15cm LH2 2 x 2 mm2 11.021 2.204 48.75 16.54 ps1,2,3=1, ps8=200; raster size set back to 2(H)x3(W) in MCC units before run 22430
  • Gmp dedicated run period

LHRS

Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Q2 [GeV2] Comment
13443 15cm LH2 2.5x3 8.841 2.503 43.0 11.89 ps1=ps2=ps3=1, ps8=20000; beam current at 60 μA; beam energy measurements started before the end of the run with tune beam. Needs to cut this period
13446, 13449 - 13453, 13455 - 13469, 13471 - 13472, 13475, 13477 - 13480 15cm LH2 2.5x3 8.841 2.503 43.0 11.89 ps1=ps2=ps3=1, ps8=20000; beam current at 60 μA
13481 15cm LH2 2.5x3 8.841 2.503 43.0 11.89 Raster failed at the end of run
13485 - 13490 15cm LH2 2.5x3 8.841 2.503 43.0 11.89 Non-standard raster setting @ 5x4, raster A X failed
13495 - 13500 15cm LH2 2.5x3 8.841 2.503 43.0 11.89 40 μA beam; Non-standard raster setting @ 5x4, raster A X failed; raster B Y failed at the end of run 13500
13504 - 13516 15cm LH2 2.5x3 8.841 2.503 43.0 11.89 40 μA beam; Raster issue fixed
13519 - 13530 15cm LH2 2.5x3 8.841 2.503 43.0 11.89 40 μA beam; Raster issue fixed

RHRS (Spectrometer position is the same as in run 22459)

Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Q2 [GeV2] Comment
22475, 22477 - 22481, 22483 - 22497, 22499 - 22500, 22502 - 22506 15cm LH2 2.5x3 8.841 2.101 48.75 12.65 ps1=ps2=ps3=1, ps8=200; beam current at 60 μA
22507 15cm LH2 2.5x3 8.841 2.101 48.75 12.65 Raster failed at the end of run
22511 - 22516 15cm LH2 2.5x3 8.841 2.101 48.75 12.65 Non-standard raster setting @ 5x4, raster A X failed
22521 - 22526 15cm LH2 2.5x3 8.841 2.101 48.75 12.65 40 μA beam; Non-standard raster setting @ 5x4, raster A X failed; raster B Y failed at the end of run 22526
22528 - 22539 15cm LH2 2.5x3 8.841 2.101 48.75 12.65 40 μA beam; Raster issue fixed
22542 - 22553 15cm LH2 2.5x3 8.841 2.101 48.75 12.65 40 μA beam

Runs on dummy target

Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Comment
21916 15cm dummy 2 x 2 mm2 4.481 1.547 52.91 Only one set of raster coils is in use; ps1,2,3=1, ps8=100;short dummy run at 10 uA; hydrogen elastic
21917 15cm dummy 2 x 2 mm2 4.481 1.547 52.91 Only one set of raster coils is in use; ps1,2,3=1, ps8=100;dummy run at 15 uA, 20 minutes no beam middle of run; hydrogen elastic
21936 15cm dummy 2 x 2 mm2 4.481 0.85 52.91 Only one set of raster coils is in use; ps1,2,3=1, ps8=100; inelastic kinematics
21937 15cm dummy 2 x 2 mm2 4.481 0.85 52.91 Only one set of raster coils is in use; ps1,2,3=1, ps8=100; inelastic kinematics
22302 15cm dummy 2 x 2 mm2 8.841 2.101 48.89 Proton elastic kinematics, raster size 2x2 mm2 in MCC unit (it was 4x2 mm2 in MCC unit which we take as 2x2 mm2 in Hall A unit)
13454 15cm dummy 2.5 x 3 mm2 8.841 2.503 43.0 Dummy run at Gmp dedicated time; beam current at 40 μA
22482 15cm dummy 2.5 x 3 mm2 8.841 2.101 48.75 Dummy run at Gmp dedicated time; beam current at 40 μA
13517, 13518 15cm dummy 2.5 x 3 mm2 8.841 2.503 43.0 Dummy run at Gmp dedicated time; beam current at 40 μA
22540, 22541 15cm dummy 2.5 x 3 mm2 8.841 2.101 48.75 Dummy run at Gmp dedicated time; beam current at 40 μA

Spectrometer pointing runs

Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Comment
21928 - 21929 Multi-foil carbon Off 4.48 0.85 52.91 ps1,2,3=1, ps8=100
12650 - 12651 Multi-foil carbon Off 4.48 1.485 37.14 DVCS config, ps9=1
22110 Multi-foil carbon Off 8.841 0.85 48.75 ps1,2,3=1, ps8=100
22173 Multi-foil carbon Off 8.841 2.101 48.75 ps1,2,3=1, ps8=200
22297 Multi-foil carbon Off 8.841 0.85 48.89 ps1,2,3=1, ps8=200
22328 Multi-foil carbon Off 11.021 0.85 48.75 ps1,2,3=1, ps8=200
22454 Multi-foil carbon Off 11.021 2.204 48.75 ps1,2,3=1, ps8=200; same spectrometer position as in run 22328

Inelastic runs on LH2

Run number Target Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Comment
21938 - 21939 15cm LH2 2 x 2 mm2 4.48 0.85 52.91 ps1,2,3=1, ps8=100

Optics commissioning runs

  • A full description of the run plan can be found here.
  • Survey results of spectrometer positions can be found here.
  • Step R1: Procedure with Sieve-slit
Run number Target Sieve Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Comment
21950 - 21967, 21969 - 21972 Multi-foil carbon Yes Off 2.30 0.75 48.75 ps1,2,3=1; Beam position not stable for runs 21950,21951; Straw chamber missing signals restored before run 21950
21973 - 21974 Single-foil carbon Yes Off 2.30 0.75 48.75 ps1,2,3=1
21988, 21990 - 21991 Single-foil carbon Yes Off 2.30 0.75 48.75 ps1,2=1,ps3=0; Noise in Cherenkov spectrum and VDC, straw chamber hit patterns is seen when triggering on s2m&&GC
21992-21995 15cm LH2 Yes 2x2 mm2 2.30 0.75 48.75 ps1,2=1, ps3=0; Noise in Cherenkov spectrum and VDC, straw chamber hit patterns is seen when triggering on s2m&&GC
  • Step R2: Delta Scan with Proton Elastic (without sieve slit)
Run number Target Sieve Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Comment
22006 15cm LH2 No 2 x 2 mm2 2.30 1.304 48.75 Delta scan ΔP=-4%; ps1,2=1
22007 15cm LH2 No 2 x 2 mm2 2.30 1.279 48.75 Delta scan ΔP=-2%; ps1,2=1
22008 15cm LH2 No 2 x 2 mm2 2.30 1.254 48.75 Delta scan ΔP=0%; ps1,2=1
22011 15cm LH2 No 2 x 2 mm2 2.30 1.229 48.75 Delta scan ΔP=+2%; statistics a bit low; ps1,2=1
22013 15cm LH2 No 2 x 2 mm2 2.30 1.203 48.75 Delta scan ΔP=+4%; ps1,2=1; dipole not stable in run 22012
  • Step R3: Dummy Target Runs
Run number Target Sieve Raster Ebeam [GeV] P0 [GeV/c] θe [deg] Comment
22009 - 22010 15cm dummy No 2 x 2 mm2 2.30 1.254 48.75 ps1,2=1

Target boiling study

  • The run plan and spectrometer settings for target current scan can be found here.
  • Carbon Target Current Scan
Run number Beam current [μA] Raster Comment
13430, 22464 60 2.5x3 LHRS: ps1=10; RHRS: ps1=6
13431, 22465 45 2.5x3 LHRS: ps1=4; RHRS: ps1=2
13433, 22466 25 2.5x3 LHRS: ps1=2; RHRS: ps1=1
13434, 22467 10 2.5x3 LHRS: ps1=1; RHRS: ps1=1
  • 15-cm Aluminum dummy
Run number Beam current [μA] Raster Comment
13435, 22468 40 2.5x3 LHRS: ps1=4; RHRS: ps1=2
  • LH2 current scan
Run number Beam current [μA] Raster Comment
13436, 22469 10 2.5x3 LHRS: ps1=5; RHRS: ps1=3
13437, 22470 20 2.5x3 LHRS: ps1=6; RHRS: ps1=3
13438, 22471 30 2.5x3 LHRS: ps1=6; RHRS: ps1=4
13439, 22472 40 2.5x3 LHRS: ps1=7; RHRS: ps1=4
13440, 22473 50 2.5x3 LHRS: ps1=8; RHRS: ps1=5
13441, 22474 60 2.5x3 LHRS: ps1=9; RHRS: ps1=6

Beam energy measurement

  • This is a collection of halog entries on the beam energy measurement. It might not be complete, and some of them might not be useful. I just keep them here for future reference.
1st Pass Bdl Measurement (beam not yet to the hall)
Bdl for flat orbit through ARC
QuickPic - harps C18A/C18B for energy measurement
NMR after restoration
Wien Angle For Hall A @ 8.8 GeV
Hall A 8.843 GeV ARC Central Ray (Bdl & NMR)
9th Dipole Bdl / NMR / Current (central ray energy 8841 MeV)
Current Beam Energy In Hall A ~ 8845 MeV (by different techniques)
Dispersive Orbit (8842 MeV)
Minor Bdl Change
Larger Wien Angle Consistent with ARC Energy Measurement Results
4th Pass Bdl & Current Check Before I vs. Bdl Scan
Hall A 9th Dipole Map (downward)
9th Dipole Mapping Issues | NMR & Position Working
9th Dipole Mapping Issues | Follow-up
Follow-up Re: Hall A 9th Dipole Map (downward)
9th Diple Bdl Scan (upward)
~11 GeV Bdl Measurement
9th Dipole Scan (upward, 100 A steps)
9th Dipole 532.7A Bdl
Finished 9th Dipole Scans
Spin precession for Hall A @ 5-pass
ARC1 & 2 Measured Beam Energy After Long Down
11GeV 5th Pass Hall A ARC Energy (Bdl only)
Monitoring ARC NMR from "cold start" of the magnet
11 GeV Dispersive Orbit !!
Shift in Bdl
2nd Pass NMR & Bdl (~4486 MeV ideal orbit)
Hall A ARC Bdl Mapper Checked
Dispersive Optics Bdl (8.8 GeV)
Hall A Energy 8842 MeV
Follow-up Re: Beam energy vs Beam current for BCM calibration run

Some cosmic runs

Run number Comment
22024 - 22025, 22065 ps1,2,3=1, ps8=200
22066 Beam back to the hall in the last four hours of the run; ps1,2,3=1, ps8=200
22073 ps1,2,3=1, ps8=200
22074 Beam back to the hall in the last one and a half hours of the run; ps1,2,3=1, ps8=200